# Local Thermal Operations and Classical Communication

### v2

@@ 7 Summary and Outlook
Finally, we consider application of LTOCC protocols to specific scenarios, starting with the task of approximating standard logical operations, CNOT and SWAP, while maintaining local equilibria of the systems under consideration. This serves to demonstrate the limitations of such operations – even when restricted to the energy-incoherent regime, LTOCC impose limits on how accurately one can perform logical gates. The second example considers possibility of generating correlations using LTOCC and their symmetric variant, demonstrating that under certain conditions it is possible to generate perfect correlations using either of the protocols.
The results presented here lay the foundation for a new framework that combines two established concepts: the distant laboratories setting and thermodynamic limitations on operations. Given the novelty of this framework, several key questions remain open for future research. The first is a deeper exploration of multi-round scenarios, which may allow for a broader range of accessible state transitions, and formulation of stronger relations between protocols with different numbers of rounds. Another important question is the introduction of coherences into the framework. This is a challenging problem, as coherent thermal operations are difficult to describe in closed form even for a single qubit [67, 43]. The distinction between thermal operations and Gibbs-preserving operations, which emerge on the quantum level, further complicates this analysis. However, taking into account the conjecture that LTOCC with shared randomness and SLTO are equal when restricted to energy-incoherent processes, extension to coherent states in future work may provide tools necessary to identify the gap between classical and genuinely quantum effects in thermodynamics of finite-size systems interacting with multiple thermal baths.
@@
Finally, the introduced framework could be applied to quantum communication protocols, potentially revealing new thermodynamic trade-offs, such as those between system temperature and achievable key rates or thermodynamical entropy of states and informational entropy of shared randomness. Such analysis would allow for a better understanding of thermally optimal operations in the context of protocols such as entanglement generation, cooling or work extraction. However, this would require studying measurements that are not restricted to the energy eigenbasis, raising questions about the inherent thermodynamic costs. These trade-offs between information-theoretic gains, like key rates, and thermodynamic costs, such as work required, represent another avenue for future research.
@@
Acknowledgments
The authors thank A. de Oliveira Jr., Jeongrak Son, Kamil Korzekwa and Karol Życzkowski for useful discussions and comments concerning the paper. Furthermore, we thank A. de Oliveira Jr. for kindly providing us with the artistic impression in Fig. 1.
RB acknowledges support by the National Science Centre, Poland, under the contract number 2021/03/Y/ST2/00193 within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733. JCz is supported by the start-up grant of the Nanyang Assistant Professorship at the Nanyang Technological University in Singapore, awarded to Nelly Ng.
@@ Appendix A Cooling maps
$\mathbf{\gamma}\approx|E_{0}\rangle\langle E_{0}|\leavevmode\nobreak\ .$ (45)
@@ Appendix B Alternative derivation of cooling maps
In the zero temperature limit $\beta\to\infty$, assuming nondegenerated Hamiltonian spectrum, the stochastic matrix $\Lambda$ corresponds to thermal operation if and only if $\Lambda_{i,j}=0$ for $i>j$. Furthermore, the extremal operations correspond to matrices with 00 and $1$ entries, restricted only by the condition $\Lambda_{ij}=0$ if $i>j$.
@@
$=\sum_{i}\quantity({\gamma}_{i}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})-\sum_{i\leq m}\quantity({\gamma}_{i}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})=1-\sum_{j>m}{\gamma}_{j}-\sum_{i\leq m}{\gamma}_{i}+\sum_{i\leq m}\sum_{j<m}\Lambda_{ij}{\gamma}_{j}$
$=\sum_{i\leq m}\sum_{j<m}\Lambda_{ij}{\gamma}_{j}\leavevmode\nobreak\ .$
@@ Appendix C Detailed discussion of one- and two-round LTOCC
Measurement: $\rho\otimes\sigma$ $\mapsto\quantity(\sum_{i}\text{Tr}(K_{i}\rho K_{i}^{\dagger})K_{i}K_{i}^{\dagger})\otimes\sigma$ (50a)
@@
Postprocessing: $\mapsto\sum_{i}p_{i}\Lambda^{(1)}(\outerproduct{E_{i}}{E_{i}})\otimes\Xi^{(1,i)}\quantity(\sigma)\otimes\outerproduct{i}{i}_{m}$ (50e)
Memory erasure: $\mapsto\sum_{i}p_{i}\Lambda^{(1)}(\outerproduct{E_{i}}{E_{i}})\otimes\Xi^{(1,i)}\quantity(\sigma)$ (50f)
where we set measurement operators $K_{i}=\quantity(\ket{E_{i}}\otimes\ket{i}_{m})\bra{E_{i}}$ | to reflect the energy-incoherent projective measurement. The Krauss operators $\tilde{K}_{ij}$ correspond to Bob’s thermal operations $\Xi^{(1,i)}$.
@@
$\mathbf{r}_{\text{LTOCC}_{1}}=\sum_{i=1}^{d}p^{(0)}_{i}\Lambda^{(1)}(\mathbf{s}(i))\otimes\Xi^{(1,i)}(\mathbf{q}^{(0)})=\sum_{i=1}^{d}p_{i}\mathbf{p}^{(1,i)}\otimes\mathbf{q}^{(1,i)}$ (51)
@@
which are depicted in center and rightmost panels of fig. 10. Note that Alice’s marginal has not been affected by the conditional operation, and thus $\sum_{ij}r_{ij}\log(\sum_{j}r_{ij})=\text{const.}$. This makes mutual information and entropy conditioned on Bob’s system equal up to an additive constant and a sign. By the same token, a similar connection exists between the entropy of the state $S(\mathbf{r})$ and the entropy conditioned on Alice’s system.
@@
The rest of the protocol proceeds in the same manner as before. Note that it is easy to reason heuristically why Corollary 3 works – conditioned processing of Bob’s state is done using local thermal operations, and thus from his perspective, it obeys the thermomajorization $\prec_{\gamma^{(B)}}$. In the post-processing step Alice uses only local thermal operation, and thus clearly obeys thermomajorization $\prec_{\gamma^{(A)}}$ locally. Taking this together with the fact that baths are not correlated, it is reasonable to expect that the joint evolution will obey thermomajorization with respect to $\prec_{\gamma^{(A)}\otimes\gamma^{(B)}}$. This is only signalised here, and actual proofs will follow in D.
@@ C.1 Remarks on relation with resource non-generating operations
$\mathcal{F}^{(A)}\otimes_{\text{max}}\mathcal{F}^{(B)}$ =$=\quantity{\rho^{(AB)}|\text{Tr}_{A}\rho^{(AB)}\in\mathcal{F}^{(B)},\,\text{Tr}_{B}\rho^{(AB)}\in\mathcal{F}^{(A)}}.$ (64)
@@
[…]
@@
$\Lambda^{(AB)}(\rho^{(AB)})=\text{Tr}_{\mathcal{B}^{(A)},\mathcal{B}^{(B)}}\left[U(\gamma^{\mathcal{B}^{(A)}}\otimes\gamma^{\mathcal{B}^{(B)}}\otimes\rho^{(AB)})U^{\dagger}\right]\leavevmode\nobreak\ .$
@@
$\Delta_{1}^{\mathcal{B}^{(A)}}=\frac{-1}{\beta^{(A)}}\log\left(\frac{\alpha}{Z_{{\mathcal{B}_{1}^{(A)}}}}\right)\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Delta_{2}^{\mathcal{B}^{(A)}}=\frac{-1}{\beta^{(A)}}\log\left(\frac{1-\alpha}{Z_{{\mathcal{B}_{2}^{(A)}}}}\right)$ (68)
@@
$\Lambda^{(AB)}(\rho^{(AB)}):=\text{Tr}_{\mathcal{B}^{(A)},\mathcal{B}^{(B)}}\left[U\left(\left((\alpha\,\gamma_{1}^{\mathcal{B}^{(A)}})\oplus((1-\alpha)\gamma_{2}^{\mathcal{B}^{(A)}})\right)\otimes(\gamma_{1}^{\mathcal{B}^{(B)}}\otimes\gamma_{2}^{\mathcal{B}^{(B)}})\otimes\rho^{(AB)}\right)U^{\dagger}\right]=$
@@
One-round Local Thermal Operations and Classical communication (LTOCC1) (with projective energy-incoherent measurements) form a subset of semilocal thermal operations (SLTO), even when considering non energy incoherent states.
@@
To show that conditioned operation $\sum_{k=1}^{n}\Lambda_{k}^{(B)}(P_{k}^{(A)}\rho^{(AB)}P_{k}^{(A)})$, where $P_{k}^{(A)}$ are projectors constituting Alice’s measurement, can be performed as SLTO, let us define a joint Bob’s bath $\mathcal{B}^{(B)}=\bigotimes_{k}\mathcal{B}_{k}^{(B)}$ with total Hamiltonian $H^{\mathcal{B}^{(B)}}=\sum_{k}\mathcal{H}_{k}^{B^{(B)}}$. We start the proof by focusing on an $n$-dimensional degenerate subspace of Alice’s bath, thus setting $\mathcal{B}^{(A)}=\mathbb{C}^{n}$ with Hamiltonian $H^{\mathcal{B}^{(A)}}\propto\mathbb{I}_{n}$, and later discuss how to generalize it to the full picture. Note that the Alice bath is completely degenerate, hence, it cannot be used to move any population between energy levels. Thus as such, it cannot be used to perform thermal operations on Alice’s subsystem. However, it may serve the role of an ancilla for energy incoherent measurements, which we are going to leverage. The global unitary in such a case is given by
@@
$=\text{Tr}_{\mathcal{B}^{(A)}\mathcal{B}^{(B)}}\Big{[}\left(\sum_{k,l}P_{k}^{(A)}\otimes U_{k}\otimes\mathbb{I}_{\mathcal{B}^{(B)}\setminus\mathcal{B}_{k}^{(B)}}\otimes|(k+l)\text{ mod }n\rangle\langle l|_{\mathcal{B}^{(A)}}\right)(\gamma^{\mathcal{B}^{(B)}}\otimes\rho^{(AB)}\otimes\frac{1}{n}\mathbb{I}_{B^{(A)}})$
@@
$=\sum_{k,l}\frac{1}{n}\text{Tr}_{\mathcal{B}^{(B)}}\left[\left(P_{k}^{(A)}\otimes U_{k}\otimes\mathbb{I}_{\mathcal{B}^{(B)}\setminus\mathcal{B}_{k}^{(B)}}\right)(\gamma^{\mathcal{B}^{(B)}}\otimes\rho^{(AB)}))\left(P_{k}^{(A)}\otimes U_{k}\otimes\mathbb{I}_{\mathcal{B}^{(B)}\setminus\mathcal{B}_{k}^{(B)}}\right)^{\dagger}\right]=$
@@
$\sum_{k}P_{k}^{(A)}\otimes\mathbb{I}_{S^{(B)}\mathcal{B}^{(B)}}\otimes\mathbb{I}_{\mathcal{B}^{(A)}}=\mathbb{I}\leavevmode\nobreak\ ,$
@@
where we used the tact that $H^{\mathcal{B}^{(A)}}$ is proportional to identity, then omitted the subspaces on which both Unitary and Hamiltonias are identity, and in the final step, we used the fact that $U_{k}$ commutes with $H^{(B)}+H_{k}^{\mathcal{B}^{(B)}}$ and $P_{k}^{(A)}$ are energy incoherent. The second commutation relation is satisfied by the same token.
@@
Combining Lemma  1 with Theorem 4 and Theorem 5, the result is obtained immediately. ∎
@@
Note that in the following proof of Theorem 7, we rely on the orthogonality of projective measurements. However, when considering energy-incoherent POVMs, the inclusion of LTOCC in SLTO is an open problem:

### v4

@@ VII Summary and Outlook
Finally, we consider the question of nonlocal correlations under LTOCC, natural in the context of the distant laboratories paradigm. We first show ability of LTOCC to generate correlations when input states are energy-incoherent, highlighting significant capability when extended by memory. Then we move on to the Bell nonlocality in the context of thermally restricted operations, for which we show a clear gap between thermally-restricted and maximal breaking of CHSH inequality. This serves to demonstrate possibility of certifying resource utility in the context of Bell nonlocality.
The results presented here lay the foundation for a new framework that combines two established concepts: the distant laboratories setting and thermodynamic limitations on operations. Given the novelty of this framework, several key questions remain open for future research. The first is a deeper exploration of multi-round scenarios, which may allow for a broader range of accessible state transitions, and formulation of stronger relations between protocols with different numbers of rounds. Next topic concerns characterization of LTOCC++M operations, which may distort Gibbs state. Contrary to LTOCC without memory, these protocols cannot be implemented without persistent classical registers, raising the question about the interplay between Landauer cost of erasing classical memory and athermality of quantum system created by corresponding operations.
Another important question is the introduction of coherences into the framework, with first steps presented in an example of detection gap for CHSH inequality. This is in general a challenging problem, as coherent thermal operations are difficult to describe in closed form even for a single qubit [74, 44]. The distinction between thermal operations and Gibbs-preserving operations, which emerge on the quantum level, further complicates this analysis. However, taking into account the conjecture that LTOCC with shared randomness and SLTO are equal when restricted to energy-incoherent processes, extension to coherent states in future work may provide tools necessary to identify the gap between classical and genuinely quantum effects in thermodynamics of finite-size systems interacting with multiple thermal baths.
@@
Finally, the introduced framework could be applied to quantum communication protocols, potentially revealing new thermodynamic trade-offs, such as those between system temperature and achievable key rates or thermodynamical entropy of states and informational entropy of shared randomness. Such analysis would allow for a better understanding of thermally optimal operations in the context of protocols such as entanglement generation, cooling or work extraction. However, this would require either shift to multicopy scenarios, similar to the one presented for the CHSH inequality breaking, or studying measurements that are not restricted to the energy eigenbasis, raising questions about the inherent thermodynamic costs. These trade-offs between information-theoretic gains, like key rates, and thermodynamic costs, such as work required, represent another avenue for future research.
@@
Acknowledgements.
The authors thank A. de Oliveira Jr., Jeongrak Son, Kamil Korzekwa and Karol Życzkowski for useful discussions and comments concerning the paper. J. Cz. would like to thank Ray Ganardi for useful comments concerning Bell nonlocality in the presence of thermal operations. Furthermore, we thank A. de Oliveira Jr. for kindly providing us with the artistic impression in Fig. 1. RB acknowledges support by the National Science Centre, Poland, under the contract number 2021/03/Y/ST2/00193 within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733. JCz is supported by the start-up grant of the Nanyang Assistant Professorship at the Nanyang Technological University in Singapore, awarded to Nelly Ng.
@@ Appendix A Cooling maps
$\mathbf{\gamma}\approx|E_{0}\rangle\langle E_{0}|\penalty 10000\ .$ (51)
@@ Appendix B Alternative derivation of cooling maps
In the zero temperature limit $\beta\to\infty$, assuming nondegenerated Hamiltonian spectrum, the stochastic matrix $\Lambda$ corresponds to thermal operation if and only if $\Lambda_{i,j}=0$ for $i>j$. Furthermore, the extremal operations correspond to matrices with 0 and $1$ entries, restricted only by the condition $\Lambda_{ij}=0$ if $i>j$.
@@
$=\sum_{i}\quantity({\gamma}_{i}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})-\sum_{i\leq m}\quantity({\gamma}_{i}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})=$
$1-\sum_{j>m}{\gamma}_{j}-\sum_{i\leq m}{\gamma}_{i}+\sum_{i\leq m}\sum_{j<m}\Lambda_{ij}{\gamma}_{j}$
$=\sum_{i\leq m}\sum_{j<m}\Lambda_{ij}{\gamma}_{j}\penalty 10000\ .$
@@ Appendix C Detailed discussion of one- and two-round LTOCC
$\text{Measurement:}\;\rho\otimes\sigma\mapsto\quantity(\sum_{i}\text{Tr}(K_{i}\rho K_{i}^{\dagger})K_{i}K_{i}^{\dagger})\otimes\sigma$ (56a)
@@
[…]
@@
$\mathbf{r}_{\text{LTOCC}_{1}}$ $=\sum_{i=1}^{d}p^{(0)}_{i}\Lambda^{(1)}(\mathbf{s}(i))\otimes\Xi^{(1,i)}(\mathbf{q}^{(0)})$ (57)
$=\sum_{i=1}^{d}p_{i}\mathbf{p}^{(1,i)}\otimes\mathbf{q}^{(1,i)}$
@@
which are depicted in center and rightmost panels of fig. 11. Note that Alice’s marginal has not been affected by the conditional operation, and thus $\sum_{ij}r_{ij}\log(\sum_{j}r_{ij})=\text{const.}$ This makes mutual information and entropy conditioned on Bob’s system equal up to an additive constant and a sign. By the same token, a similar connection exists between the entropy of the state $S(\mathbf{r})$ and the entropy conditioned on Alice’s system.
@@
The rest of the protocol proceeds in the same manner as before. Note that it is easy to reason heuristically why Corollary 3 works – conditioned processing of Bob’s state is done using local thermal operations, and thus from his perspective, it obeys the thermomajorization $\prec_{\gamma^{(B)}}$. In the post-processing step Alice uses only local thermal operation, and thus clearly obeys thermomajorization $\prec_{\gamma^{(A)}}$ locally. Taking this together with the fact that baths are not correlated, it is reasonable to expect that the joint evolution will obey thermomajorization with respect to $\prec_{\gamma^{(A)}\otimes\gamma^{(B)}}$. This is only signalised here, and actual proofs will follow in Appendix D.
@@ C.1 Remarks on relation with resource non-generating operations
$\mathcal{F}^{(A)}\otimes_{\text{max}}\mathcal{F}^{(B)}$ $=\quantity{\rho^{(AB)}|\text{Tr}_{A}\rho^{(AB)}\in\mathcal{F}^{(B)},\,\text{Tr}_{B}\rho^{(AB)}\in\mathcal{F}^{(A)}}$. (70)
@@
$\text{LTOCC+R}\subset\,\text{RNG}\quantity(\mathcal{F}^{(A)}_{\text{TO}}\otimes_{\text{min}}\mathcal{F}^{(B)}_{\text{TO}})$ (73)
@@
$\Lambda^{(AB)}(\rho^{(AB)})=\text{Tr}_{\mathcal{B}^{(A)},\mathcal{B}^{(B)}}\left[U(\gamma^{\mathcal{B}^{(A)}}\otimes\gamma^{\mathcal{B}^{(B)}}\otimes\rho^{(AB)})U^{\dagger}\right]\penalty 10000\ .$
@@
$\Delta_{1}^{\mathcal{B}^{(A)}}=\frac{-1}{\beta^{(A)}}\log\left(\frac{\alpha}{Z_{{\mathcal{B}_{1}^{(A)}}}}\right)\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \Delta_{2}^{\mathcal{B}^{(A)}}=\frac{-1}{\beta^{(A)}}\log\left(\frac{1-\alpha}{Z_{{\mathcal{B}_{2}^{(A)}}}}\right)$ (74)
@@
$\Lambda^{(AB)}(\rho^{(AB)})=\text{Tr}_{\mathcal{B}^{(A)},\mathcal{B}^{(B)}}\left[U\left(\left((\alpha\,\gamma_{1}^{\mathcal{B}^{(A)}})\oplus((1-\alpha)\gamma_{2}^{\mathcal{B}^{(A)}})\right)\otimes(\gamma_{1}^{\mathcal{B}^{(B)}}\otimes\gamma_{2}^{\mathcal{B}^{(B)}})\otimes\rho^{(AB)}\right)U^{\dagger}\right]=$
@@
One-round Local Thermal Operations and Classical communication (LTOCC1) (with projective energy-incoherent measurements) form a subset of semilocal thermal operations (SLTO), even when considering non energy-incoherent states.
@@
To show that conditioned operation $\sum_{k=1}^{n}\Lambda_{k}^{(B)}(P_{k}^{(A)}\rho^{(AB)}P_{k}^{(A)})$, where $P_{k}^{(A)}$ are projectors constituting Alice’s measurement, can be performed as SLTO, let us define a joint Bob’s bath $\mathcal{B}^{(B)}=\bigotimes_{k}\mathcal{B}_{k}^{(B)}$ with total Hamiltonian $H^{\mathcal{B}^{(B)}}=\sum_{k}\mathcal{H}_{k}^{B^{(B)}}$. We start the proof by focusing on an $n$-dimensional degenerate subspace of Alice’s bath, thus setting $\mathcal{B}^{(A)}=\mathbb{C}^{n}$ with Hamiltonian $H^{\mathcal{B}^{(A)}}\propto\mathbb{I}_{n}$, and later discuss how to generalize it to the full picture. Note that the Alice bath is completely degenerate, hence, it cannot be used to move any population between energy levels. Thus as such, it cannot be used to perform thermal operations on Alice’s subsystem. However, it may serve the role of an ancilla for energy-incoherent measurements, which we are going to leverage. The global unitary in such a case is given by
@@
$=\text{Tr}_{\mathcal{B}^{(A)}\mathcal{B}^{(B)}}\Big[\left(\sum_{k,l}P_{k}^{(A)}\otimes U_{k}\otimes\mathbb{I}_{\mathcal{B}^{(B)}\setminus\mathcal{B}_{k}^{(B)}}\otimes|(k+l)\text{ mod }n\rangle\langle l|_{\mathcal{B}^{(A)}}\right)\left(\gamma^{\mathcal{B}^{(B)}}\otimes\rho^{(AB)}\otimes\frac{1}{n}\mathbb{I}_{B^{(A)}}\right)$
@@
$=\sum_{k,l}\frac{1}{n}\text{Tr}_{\mathcal{B}^{(B)}}\left[\left(P_{k}^{(A)}\otimes U_{k}\otimes\mathbb{I}_{\mathcal{B}^{(B)}\setminus\mathcal{B}_{k}^{(B)}}\right)(\gamma^{\mathcal{B}^{(B)}}\otimes\rho^{(AB)})\left(P_{k}^{(A)}\otimes U_{k}\otimes\mathbb{I}_{\mathcal{B}^{(B)}\setminus\mathcal{B}_{k}^{(B)}}\right)^{\dagger}\right]=$
@@
$\sum_{k}P_{k}^{(A)}\otimes\mathbb{I}_{S^{(B)}\mathcal{B}^{(B)}}\otimes\mathbb{I}_{\mathcal{B}^{(A)}}=\mathbb{I}\penalty 10000\ ,$
@@
where we used the tact that $H^{\mathcal{B}^{(A)}}$ is proportional to identity, then omitted the subspaces on which both Unitary and Hamiltonias are identity, and in the final step, we used the fact that $U_{k}$ commutes with $H^{(B)}+H_{k}^{\mathcal{B}^{(B)}}$ and $P_{k}^{(A)}$ are energy-incoherent. The second commutation relation is satisfied by the same token.
@@
Combining Lemma 2 with Theorem 4 and Theorem 5, the result is obtained immediately. ∎
@@
In the proof of Theorem 7, we rely on the orthogonality of projective measurements, in order to guarantee unitarity of joint evolution. However, even when considering energy-incoherent POVMs, which satisfy the minimal requirement of (average) Gibbs-preservation, this orthogonality ceases to hold. Thus above reasoning cannot be directly applied for beyond PVMs.
Furthermore, as presented below, the action of an energy-incoherent POVM is not equivalent to a convex combination of energy-incoherent PVM. Thus, one cannot construct LTOCC protocol utilizing POVM by probabilistic mixture of LTOCC protocols based on PVMs.
Observation 9.
Non-projective energy-incoherent POVM cannot be implemented as a convex combination of energy-incoherent PVMs.
Proof.
[…]
In this setting one can obtain the probability of measuring basis state $|E_{j}\rangle$ by POVM element $F_{i}$ in two different ways. The direct calculation gives $p=\text{Tr}[F_{i}|E_{j}\rangle\langle E_{j}|]=q_{ij}$, while the composition of PVMs yields $p=\text{Tr}[F_{i}|E_{j}\rangle\langle E_{j}|]=\sum_{k}r_{ik}\text{Tr}[P_{ik}|E_{j}\rangle\langle E_{j}|]=\sum_{k}r_{ik}d_{ik}(j)$. Thus we obtain consistency requirement $q_{ij}=\sum_{k}r_{ik}d_{ik}(j)$.
The application of POVM on arbitrary quantum state $\rho$ gives, up to normalization $p_{i}$,
$p_{i}\rho^{(i)}=\sqrt{F_{i}}\rho\sqrt{F_{i}}=\sum_{j,j^{\prime}}\sqrt{q_{ij}q_{ij^{\prime}}}\;|E_{j}\rangle\langle E_{j}|\rho|E_{j^{\prime}}\rangle\langle E_{j^{\prime}}|=\sum_{j,j^{\prime}}\sqrt{\sum_{k,k^{\prime}}r_{ik}d_{ik}(j)r_{ik^{\prime}}d_{ik^{\prime}}(j^{\prime})}\;|E_{j}\rangle\langle E_{j}|\rho|E_{j^{\prime}}\rangle\langle E_{j^{\prime}}|.$ (78)
On the other hand, the combination of projective measurements creates post-measurement state
$\sum_{k}r_{ik}P_{ik}\rho P_{ik}=\sum_{j,j^{\prime}}\sum_{k}r_{ik}d_{ik}(j)d_{ik}(j^{\prime})\;|E_{j}\rangle\langle E_{j}|\rho|E_{j^{\prime}}\rangle\langle E_{j^{\prime}}|.$ (79)
Comparing these two equations element-wise, one obtains
$\forall_{j,j^{\prime}}\;\;\sum_{k}r_{ik}d_{ik}(j)d_{ik}(j^{\prime})=\sqrt{\sum_{k,k^{\prime}}r_{ik}d_{ik}(j)r_{ik^{\prime}}d_{ik^{\prime}}(j^{\prime})},$
which, after squaring both sides, and small rearagement of terms results in
$\forall_{j,j^{\prime}}\;\;\sum_{kk^{\prime}}\Big(1-d_{ik}(j^{\prime})d_{ik^{\prime}}(j)\Big)d_{ik}(j)d_{ik^{\prime}}(j^{\prime})r_{ik}r_{ik^{\prime}}=0.$
Each term in the above sum is nonnegative, so the sum may be equal zero only if each summand is zero. Furthermore, since $r_{ik}>0$ we may equivalently write
$\forall_{j,j^{\prime}}\forall_{k,k^{\prime}}\;\;\Big(1-d_{ik}(j^{\prime})d_{ik^{\prime}}(j)\Big)d_{ik}(j)d_{ik^{\prime}}(j^{\prime})=0.$ (80)
If the measurement $F_{i}$ is not projective, then some projectors $\{P_{ik}\}_{k=0}^{k_{max}(i)}$ must differ from each other. Let $P_{ik_{0}}$ and $P_{ik_{1}}$ be two such different projectors, with $|E_{j_{0}}\rangle$ being a basis state such that $P_{ik_{0}}|E_{j_{0}}\rangle=|E_{j_{0}}\rangle$ and $P_{ik_{1}}|E_{j_{0}}\rangle=0$. Substituting $k,k^{\prime},j$ by $k_{0},k_{1},j_{0}$ the equation (80) simplifies to $\forall_{j^{\prime}}\;\;d_{ik^{\prime}}(j^{\prime})=0,$ which means that the entire projection $P_{ik_{1}}$ is equal 0. Therefore energy-incoherent POVM cannot be considered as any convex combination, or probabilistic post-processing, of an energy-incoherent PVM. ∎
The above observation can be strengthened by separately considering the diagonal and off-diagonal elements of the output state, which leads to the conclusion that POVMs in general preserve more coherence than composition of PVMs with the same measurement probabilities. This excludes the implementation of POVM as a combination of PVMs followed by decohering thermal post-processing.
With these observations in mind, together with the discrepancy between thermal operations and Gibbs-preserving operations for coherent states, we believe that the extension of Theorem 7 by POVMs, if possible, requires significantly different proof techniques. Therefore, the problem of the inclusion of LTOCC in SLTO remains open: